# Jan Blumenkamp, Ajay Shankar, Matteo Bettini, Joshua Bird, Amanda Prorok

### v1

@@ Abstract
Compact robotic platforms with powerful compute and actuation capabilities are key enablers for practical, real-world deployments of multi-agent research. This article introduces a tightly integrated hardware, control, and simulation software stack on a fleet of holonomic ground robot platforms designed with this motivation. Our robots, a fleet of customised DJI Robomaster S1 vehicles, offer a balance between small robots that do not possess sufficient compute or actuation capabilities and larger robots that are unsuitable for indoor multi-robot tests. They run a modular ROS2-based optimal estimation and control stack for full onboard autonomy, contain ad-hoc peer-to-peer communication infrastructure, and can zero-shot run multi-agent reinforcement learning (MARL) policies trained in our vectorized multi-agent simulation framework. We present an in-depth review of other platforms currently available, showcase new experimental validation of our system’s capabilities, and introduce case studies that highlight the versatility and reliabilty of our system as a testbed for a wide range of research demonstrations. Our system as well as supplementary material is available online. 111https://proroklab.github.io/cambridge-robomaster
@@ 1 Introduction
Multi-agent robotics research necessitates robotic platforms that can be used as deployment testbeds for rapid development and evaluations [32, 16, 52, 6, 41, 8, 34, 38, 54, 43, 50, 28, 42, 7, 14, 11, 13, 27, 55]. The choices for such platforms are usually informed by several factors such as the configuration space of the problem, the number of agents involved, and the computational sophistication expected from each agent. For instance, small-scale robots, such as the Turtlebot [51], will frequently be limited in their agility of motion and are restricted in their onboard computing capabilities. On the other hand, while larger robots (Jackal [18] or AutoRally [53]) have greater maneuverability and can carry more compute payloads, operating multiple of them in smaller research spaces can be cost- and space- prohibitive. Successfully testing and developing multi-robot solutions requires platforms that strike a fine balance between these objectives.
In this article, we present an agile and affordable system and platform, shown in Fig. 1 as a multi-robot team setup. Our system, an omnidirectional indoor ground robot based on the DJI Robomaster S1, is affordable (starting at USD 689 in the base configuration to USD 1633 with all sensors and an Jetson Orin NX), agile (we demonstrate indoor trajectory tracking capabilities of up to and accelerations of up to , versatile (we show experiments in classical control to zero-shot deployment of MARL policies trained in our simulator VMAS [10]), and is easily replicable (our entire stack is open-sourced, requiring minimal construction time). These have been made possible due to our retrofitting of the onboard compute module and careful reverse engineering of the main communication interface. We present two further contributions in this work. Second, through an expansive review of related systems and literature, we contrast our solution against others’ and position our system at the frontier of low-cost agile research platforms. Third, we describe enhancements to our prior work (VMAS [10]) that enable rapid deployment of control policies learnt in simulation onto this platform. We present evaluations and case studies that corroborate the claims presented above, and showcase the wide range of research avenues that this system makes accessible.
@@ 2 Related Work
In the last decade, a wide range of (multi-) robot platforms have been introduced to the research community. Here we present an in-depth literature review of commercial as well as research platforms. Our classification scheme will differentiate between them based on their dynamics (Ackermann, differential, and omnidirectional drive, as well as airborne/multirotor platforms), and also based on their specific use-cases in single- or multi-agent research. Since we are primarily interested in multi-robot research, we investigate the tradeoff between cost and speed of these platforms, as agility is an attribute we require in the team (note that we use ‘speed’ and ‘agility’ interchangeably – this is a justified simplification for indoor robots operating in confined spaces), and cost limits the size of the team.
@@
EPFL | Cellulo [42] | 0.15 | 150 | 0.18 | ✓ | ||
Lincoln Uni | Colias [7] | 0.35 | 40 | 0.01 | ✓ | ||
Hiwonder | Jetauto [46] | 0.50 | 650 | 4.00 | ✓ | ||
Hangfa | Compass Q2 [30] | 0.75 | 3,500 | 16.00 | ✓ | ||
KUKA | youBot [12] | 0.80 | 30,000 | 20.00 | ✓ | ||
Clearpath | Ridgeback [19] | 1.10 | 50,000 | 135.00 | ✓ | ||
Hangfa | Navigator R2 [31] | 1.20 | 20,300 | 35.00 | ✓ | ||
Clearpath | Dingo-O [17] | 1.30 | 10,000 | 13.00 | ✓ | ||
RoboWorks | Mecabot X [47] | 1.32 | 6,000 | 20.50 | ✓ | ||
Lynxmotion | A4WD3 [37] | 1.40 | 900 | 5.55 | ✓ | ||
RDK | X3 [44] | 1.50 | 590 | 1.93 | ✓ | ||
Agilex | Ranger Mini [2] | 1.50 | 12,400 | 64.50 | ✓ | ||
RoboWorks | Mecabot Pro [46] | 1.83 | 2,600 | 10.80 | ✓ | ||
Clearpath | Boxer [20] | 2.00 | 60,000 | 127.00 | ✓ | ||
Festo | Robotino [25] | 2.70 | 8,000 | 22.80 | ✓ | ||
Cambridge | RoboMaster | 4.45 | 670 | 3.00 | ✓ | ✓ |
Fig. 2 visualises our survey in a 2D plot, and Tab. 1 provides detailed statistics specifically for for omnidirectional drive robots. We find that the majority of multi-agent research platforms utilize differential drive [41, 34, 6, 43, 54, 38, 50, 8, 16], thus positioning them in the bottom left corner of Fig. 2. The reason is that such platforms tend to be designed to minimize cost, with 10s of such agents in mind. The slowest of these platforms is Alice [16] at , and the fastest is the Khepera IV [50] with . Three ground-based platforms are specifically geared towards multi-agent research but use dynamics other than differential drive, specifically the Minicar [32], Cellulo[42], and Colias [7]. More recently, slightly faster commercial differential drive robots have become popular [18, 21, 51, 22], ranging from for the Turtlebot 3 [51] to for the Diabolo [22].
From the bottom left, we can see two clusters emerging towards the top and the right. The omnidirectional drive robots [17, 19, 25, 12, 31, 47, 30, 46, 46, 42, 7, 44, 20, 37, 3, 2] tend to move faster than differential drive, ranging from (Cellulo [42]) to [37], but are also significantly more expensive, with a mean cost of USD 14k.
@@
Our solution sits at a frontier beyond which agility (speed) is attained only by aerial platforms, and a lower cost is only attained by sacrificing agility. Furthermore, as we will describe in later sections, our software stack supports a wide range of relevant research problems, and has a proven track-record throughout six previous publications over the last three years [14, 11, 13, 27, 55, 15].
@@ 3.1 Platform
The onboard computer interfaces with a sub-controller, managing all sensors and wheels. We replaced the onboard computer, gimbal, and blaster gun with a custom computing solution. We laser-cut components from acrylic to mount our custom expansion in place of the turret. All our design files are also part of our open-source release. 222https://github.com/proroklab/cambridge-robomaster
@@ 3.2 Compute
The stock onboard compute module on the RoboMasters is impractical for running custom algorithms or implementing additional interfaces to cameras, sensors, networks, etc. We therefore upgraded the onboard computer to a more powerful and versatile single board computer (SBC), focusing on the NVidia Jetson Orin family for its recency and lifecycle availability until January 2030333https://developer.nvidia.com/embedded/lifecycle, accessed 24/02/22. While the Jetson Orin NX 16GB was selected for its robust processing capabilities and suitability for our research, our design easily admits other cost-effective single-board-computers such as the Raspberry Pi. The Jetson Orin series necessitates a carrier board. Our choice, the AverMedia D131 carrier board, was influenced by the availability of a CAN interface – essential for communicating with the RoboMaster subcontroller – and an optimal balance of interface availability and cost. It accommodates two PCIe M.2 slots, where we installed a 256 GB SSD and an Intel 9260 WiFi module. In an alternative Raspberry Pi setup, on the other hand, we used the built-in ports and interfaces and added an additional CAN shield for communication with the RoboMaster base.
@@ 3.3 CAN Communication
DJI offers an API for the RoboMaster EP, the educational counterpart to the RoboMaster S1, which enables control via a secondary computer (e.g., a Raspberry Pi) through USB. However, this platform is not listed for sale on the DJI store, is unavailable in many parts of the world (including the UK), and at around USD 1500, is significantly costlier than the S1 model. The API, which DJI has made open-source444https://github.com/dji-sdk/RoboMaster-SDK, accessed 03/19/24, served as the foundation for our reverse engineering of the communication protocol between the main computer and the sub-controller. DJI employs Controller Area Network (CAN) for hardware layer transport but overlays it with a simplified serial package-based protocol. This design likely addresses the limitations of CAN’s 8-byte data size per packet while facilitating multi-node network connectivity, leveraging CAN’s collision and arbitration management capabilities. We outline the protocol’s structure in Fig. 4, noting its use of a higher-level transport layer to specify node source and destination IDs, sequence counters for package tracking, various command sets and IDs, and acknowledgments, enabling subscription to specific package IDs for continuous data stream over the network. We release our C++ implementation of the messaging protocol as stand-alone repository.555https://github.com/proroklab/robomaster˙sdk˙can
@@ 3.4 Infrastructure
Our system employs dual network interfaces to facilitate distinct communication channels: a backbone infrastructure network for all participating systems, and, an ad-hoc network exclusively for inter-robot communication. For infrastructure-related communication, we utilize an Intel 9260 PCIe module for the Jetson and the integrated WiFi module for the Raspberry Pi that connects to a central WiFi router (described below). This setup integrates with broader infrastructure components, such as desktop systems operating the Multi-Robot Manager UI, a motion capture system, and essential hardware like an emergency-stop button. This network also allows for SSH access into the robots, file sharing, and centralized ROS messaging. In scenarios requiring only centralized control, a desktop machine executes our entire control stack (Freyja, described later) for each robot, and transmits low-level wheel commands to each robot via ROS2.
Conversely, for direct robot-to-robot communication, we use Netgear A6210 USB3.0 modules to establish an adhoc multicast network. This configuration does not rely on external infrastructure, enabling seamless peer-to-peer network interactions.
@@
The deployment of an enterprise-grade router proved crucial for centralized robot operation. We opted for a Mikrotik Router, selected for its superior configurability and flexible antenna setup. We employ a structured IP assignment strategy within the 10.x.y.z address range across all devices, and bridge appropriate subnetworks that contain different categories of devices (robots, computers, guest-devices, etc).
Our router and ad-hoc setups typically operate at a broadcast channel capacity of , and we typically observe around network load when operating 5 robots concurrently in a centralized mode (transmitting wheel velocity commands to each robot at 50 Hz).
@@
Crash prevention is a priority in high-velocity multi-robot experiments, and can be particularly complex to implement in a fully decentralized framework. Inevitable system latencies in re-establishing SSH sessions and the finite interval between terminating a control process and robot immobilization necessitates a more reliable stop mechanism. Consequently, we integrate an emergency stop service into our RoboMaster ROS driver, accessible via the infrastructure network, that immediately halts velocity commands at the lowest level. Additionally, we developed a stand-alone physical emergency stop button, powered by a Raspberry Pi Zero. This button connects to the infrastructure network and features color LEDs to display the connectivity and status of the robots.
@@
Incorporating on-robot displays has proven advantageous, especially for operations outside the conventional laboratory setting where an infrastructure network may be absent. These displays, compact and economical at pixels, present vital statistics including CPU and memory usage, CPU temperature, network details (such as IP addresses and interface connectivity), battery status, and the robot’s hostname. An integrated button allows users to navigate through different information pages and, crucially, activates the emergency stop service if the robot is in motion. This feature is particularly useful in field deployments, offering a direct control and monitoring mechanism when network access is unavailable.
@@ 3.7 Simulation: VMAS
To enable this deployment pipeline, we integrated several new features into the simulator. Firstly, to generate the input required by Freyja, we implement a PID controller layer that regulates the input forces required by VMAS based on these reference velocities. This layer allows us to train neural network policies that output desired velocities and directly use them on our physical platform, without any change. Secondly, a key to the real-world deployment was an iterative fine-tuning process for some of the physics parameters in VMAS. These are parameters that depend on the robot platform and its interaction with the environment, such as: friction coefficients, drag, control ranges, state boundaries, and robot geometry. This tuning process can be done once for each robot-environment pair and can be then utilized for any subsequent deployments.
A model for the RoboMaster platform in VMAS enables applications beyond prototyping in simulation and deploying in the real world. For instance, thanks to the accelerated compute available on our platform, policies trained in VMAS on approximated representations of the environment can be fine-tuned on the robot using data collected in the real world.
@@ 4 Evaluations
We first demonstrate the agility of our platform in a single-robot experiment in which we benchmark the tracking performance of Freyja as well as the velocity of our robot platform at high speeds. We demonstrate this on a straight line trajectory and a circle trajectory in Fig. 6. The straight line trajectory has a length of ca. , and the circle trajectory a diameter of . We show that the robot reaches a top velocity of , thus exceeding the maximum rated speed of the DJI RoboMaster platform by , and accelerations of up to in the straight line trajectory with a maximum error of , and maintains a velocity of on the circle trajectory with an average error of . In the line trajectory, we are constrained by the size of the laboratory, which prohibits the robots from going faster.
@@
for a multi-robot navigation task trained in VMAS [10] following the pipeline described in Sec. 3.7.
Agents are rewarded for navigating to their goal while avoiding collisions with each other, to do so they use the GPPO model ([11]), which leverages a Graph Neural Network (GNN) in the policy for inter-agent communication. The GNN utilizes a 5-layer MLP with a total of 2300 trainable parameters, which can be evaluated on the Jetson Orin NX in . The results, reported in Fig. 7, show that the agents are able to seamlessly execute the learned policy in the previously unseen real-world environment.
@@
a nonlinear model predictive controller avoid collisions with both peers and static obstacles.
Fig. 8 tests a collision avoidance scenario at an intersection, where two robots (traveling along the horizontal and vertical axes) would nominally collide at the intersection.
However, they are able localize each other close to the origin when common features are available in the environment, and the robot traveling along the vertical axis is able to slow down to avoid colliding.
Multi-robot control using CoViS-Net Foundation Model.
In this experiment, we deploy our CoViS-Net foundation model [15], a pre-trained Neural Network to estimate relative poses between two camera frames, on two different robots. The camera images of two robots are encoded, and the features are broadcasted over the ad-hoc network to other robots within communication range. A second model then compares the features in the image to predict the relative pose, which is then used to track a trajectory using Freyja.
@@
Case Studies. Our system has reached this level of maturity over the course of several iterations of research and development in the last four years. Apart from the capabilities demonstrated above, this framework has featured as a primary research platforms in several prior multi-agent work from our group:

### v2

@@ Abstract
Compact robotic platforms with powerful compute and actuation capabilities are key enablers for practical, real-world deployments of multi-agent research. This article introduces a tightly integrated hardware, control, and simulation software stack on a fleet of holonomic ground robot platforms designed with this motivation. Our robots, a fleet of customised DJI Robomaster S1 vehicles, offer a balance between small robots that do not possess sufficient compute or actuation capabilities and larger robots that are unsuitable for indoor multi-robot tests. They run a modular ROS2-based optimal estimation and control stack for full onboard autonomy, contain ad-hoc peer-to-peer communication infrastructure, and can zero-shot run multi-agent reinforcement learning (MARL) policies trained in our vectorized multi-agent simulation framework. We present an in-depth review of other platforms currently available, showcase new experimental validation of our system’s capabilities, and introduce case studies that highlight the versatility and reliability of our system as a testbed for a wide range of research demonstrations. Our system as well as supplementary material is available online. 111https://proroklab.github.io/cambridge-robomaster
@@ 1 Introduction
Multi-agent robotics research necessitates robotic platforms that can be used as deployment testbeds for rapid development and evaluations [36, 18, 54, 6, 45, 8, 38, 42, 56, 46, 52, 31, 58, 7, 15, 12, 14, 30, 57]. The choices for such platforms are usually informed by several factors such as the configuration space of the problem, the number of agents involved, and the computational sophistication expected from each agent. For instance, small-scale robots, such as the Turtlebot [53], will frequently be limited in their agility of motion and are restricted in their onboard computing capabilities. On the other hand, while larger robots (Jackal [20] or AutoRally [55]) have greater manoeuvrability and can carry more compute payloads, operating multiple of them in smaller research spaces can be cost- and space- prohibitive. Successfully testing and developing multi-robot solutions requires platforms that strike a fine balance between these objectives.
In this article, we present an agile and affordable system and platform, shown in Fig. 1 as a multi-robot team setup. Our system, an omnidirectional indoor ground robot based on the DJI Robomaster S1, is affordable (starting at USD 689 in the base configuration to USD 1633 with all sensors and an Jetson Orin NX), agile (we demonstrate indoor trajectory tracking capabilities of up to and accelerations of up to ), versatile (we show experiments from classical control to zero-shot deployment of MARL policies trained in our simulator VMAS [10]), and is easily replicable (our entire stack is open-sourced, requiring minimal construction time). These have been made possible due to our retrofitting of the onboard compute module and careful reverse engineering of the main communication interface. We present two further contributions in this work. First, through an expansive review of related systems and literature, we contrast our solution against others’ and position our system at the frontier of low-cost agile research platforms. Second, we describe enhancements to our prior work (VMAS [10]) that enable rapid deployment of control policies learnt in simulation onto this platform. We present evaluations and case studies that corroborate the claims presented above, and showcase the wide range of research avenues that this system makes accessible.
@@ 2 Related Work
In the last decade, a wide range of (multi-) robot platforms have been introduced to the research community. Here we present an in-depth literature review of commercial as well as research platforms. Our classification scheme will differentiate between them based on their dynamics (Ackermann, differential, and omnidirectional drive, as well as airborne/multirotor platforms), and also based on their specific use cases in single- or multi-agent research. Since we are primarily interested in multi-robot research, we investigate the tradeoff between cost and speed of these platforms, as agility is an attribute we require in the team (note that we use ‘speed’ and ‘agility’ interchangeably – this is a justified simplification for indoor robots operating in confined spaces), and cost limits the size of the team.
@@
| EPFL | Cellulo [58] | 0.15 | 150.0 | 0.18 | ✓ | ||
| Lincoln Uni | Colias [7] | 0.35 | 40.0 | 0.01 | ✓ | ||
| Hiwonder | Jetauto [48] | 0.50 | 650.0 | 4.00 | ✓ | ||
| Hangfa | Compass Q2 [33] | 0.75 | 3,500.0 | 16.00 | ✓ | ||
| KUKA | youBot [13] | 0.80 | 30,000.0 | 20.00 | ✓ | ||
| Agilex | Limo [2] | 1.00 | 3,600.0 | 4.80 | ✓ | ||
| Clearpath | Ridgeback [21] | 1.10 | 50,000.0 | 135.00 | ✓ | ||
| Hangfa | Navigator Q2 [34] | 1.20 | 20,300.0 | 35.00 | ✓ | ||
| Clearpath | Dingo-O [19] | 1.30 | 10,000.0 | 13.00 | ✓ | ||
| RoboWorks | Mecabot X [49] | 1.32 | 6,000.0 | 20.50 | ✓ | ||
| Lynxmotion | A4WD3 [41] | 1.40 | 900.0 | 5.55 | ✓ | ||
| RDK | X3 [47] | 1.50 | 590.0 | 1.93 | ✓ | ||
| Agilex | Ranger Mini [3] | 1.50 | 12,400.0 | 64.50 | ✓ | ||
| RoboWorks | Mecabot Pro [48] | 1.83 | 2,600.0 | 10.80 | ✓ | ||
| Clearpath | Boxer [22] | 2.00 | 60,000.0 | 127.00 | ✓ | ||
| Festo | Robotino [28] | 2.70 | 8,000.0 | 22.80 | ✓ | ||
| Cambridge | RoboMaster (ours) | 4.45 | 689.0 | 3.00 | ✓ | ✓ |
Fig. 2 visualises our survey in a 2D plot, and Tab. 1 provides detailed statistics specifically for for omnidirectional drive robots. We find that the majority of multi-agent research platforms utilize differential drive [45, 38, 6, 46, 56, 42, 52, 8, 18], thus positioning them in the bottom left corner of Fig. 2. The reason is that such platforms tend to be designed to minimize cost, with tens of such agents in mind. The slowest of these platforms is Alice [18] at , and the fastest is the Khepera IV [52] with . Three ground-based platforms are specifically geared towards multi-agent research but use dynamics other than differential drive, specifically the Minicar [36], Cellulo[58], and Colias [7]. More recently, slightly faster commercial differential drive robots have become popular [20, 23, 53, 25, 35], ranging from for the Turtlebot 3 [53] to for the Diabolo [25].
From the bottom left, we can see two clusters emerging towards the top and the right. The omnidirectional drive robots [58, 7, 48, 33, 13, 21, 34, 19, 49, 47, 3, 48, 22, 28, 41, 2] tend to move faster than differential drive, ranging from (Cellulo [58]) to (Festo Robotino [28]), but are also significantly more expensive, with a mean cost of USD 13k.
@@
Our solution sits at a frontier beyond which agility (speed) is attained only by aerial platforms, and a lower cost is only attained by sacrificing agility. Furthermore, as we will describe in later sections, our software stack supports a wide range of relevant research problems, and has a proven track record throughout six previous publications over the last three years [15, 12, 14, 30, 57, 16].
@@ 3.1 Platform
The onboard computer interfaces with a sub-controller, managing all sensors and wheels. We replaced the onboard computer, gimbal, and blaster gun with a custom computing solution. We laser-cut components from acrylic to mount our custom expansion in place of the turret. We provide a detailed overview of the compute configurations in Fig. 3. All our design files are also part of our open-source release. 222https://github.com/proroklab/cambridge-robomaster
@@ 3.2 Compute
The stock onboard compute module on the RoboMasters is impractical for running custom algorithms or implementing additional interfaces to cameras, sensors, networks, etc. We therefore upgraded the onboard computer to a more powerful and versatile single board computer (SBC), focusing on the NVidia Jetson Orin family for its recency and lifecycle availability until January 2030333https://developer.nvidia.com/embedded/lifecycle, accessed 24/02/22. While the Jetson Orin NX 16GB was selected for its robust processing capabilities and suitability for our research, our design easily admits other cost-effective single-board computers such as the Raspberry Pi. The Jetson Orin series necessitates a carrier board. Our choice, the AverMedia D131 carrier board, was influenced by the availability of a CAN interface – essential for communicating with the RoboMaster sub-controller – and an optimal balance of interface availability and cost. It accommodates two PCIe M.2 slots, where we installed a 256 GB SSD and an Intel 9260 WiFi module. In an alternative Raspberry Pi setup, we used the built-in ports and interfaces and added an additional CAN shield for communication with the RoboMaster base. We provide a detailed bill of material in Tab. 2.
@@ 3.3 CAN Communication
DJI offers an API for the RoboMaster EP, the educational counterpart to the RoboMaster S1, which enables control via a secondary computer (e.g., a Raspberry Pi) through USB. However, this platform is not listed for sale on the DJI store, is unavailable in many parts of the world (including the UK), and at around USD 1500, is significantly costlier than the S1 model. The API, which DJI has made open-source444https://github.com/dji-sdk/RoboMaster-SDK, accessed 03/19/24, served as the foundation for our reverse engineering of the communication protocol between the main computer and the sub-controller. DJI employs a Controller Area Network (CAN) for hardware layer transport but overlays it with a simplified serial package-based protocol. This design likely addresses the limitations of CAN’s 8-byte data size per packet while facilitating multi-node network connectivity, leveraging CAN’s collision and arbitration management capabilities. We outline the protocol’s structure in Fig. 4, noting its use of a higher-level transport layer to specify node source and destination IDs, sequence counters for package tracking, various command sets and IDs, and acknowledgements, enabling subscription to specific package IDs for continuous data stream over the network. We release our C++ implementation of the messaging protocol as a stand-alone repository.555https://github.com/proroklab/robomaster˙sdk˙can
@@ 3.4 Infrastructure
Our system employs dual network interfaces to facilitate distinct communication channels: a backbone infrastructure network for all participating systems, and, an ad-hoc network exclusively for inter-robot communication. For infrastructure-related communication, we utilize an Intel 9260 PCIe module for the Jetson and the integrated WiFi module for the Raspberry Pi that connects to a central WiFi router (described below). This setup integrates with broader infrastructure components, such as desktop systems operating the Multi-Robot Manager UI, a motion capture system, and essential hardware like an emergency stop button. This network also allows for SSH access into the robots, file sharing, and centralized ROS messaging. In scenarios requiring only centralized control, a desktop machine executes our entire control stack (Freyja, described later) for each robot, and transmits low-level wheel commands to each robot via ROS2.
For direct robot-to-robot communication, we use Netgear A6210 USB3.0 modules to establish an adhoc multicast network. This configuration does not rely on external infrastructure, enabling seamless peer-to-peer interactions.
@@
The deployment of an enterprise-grade router proved crucial for centralized robot operation. We opted for a Mikrotik Router, selected for its superior configurability and flexible antenna setup. We employ a structured IP assignment strategy and bridge appropriate subnetworks that contain different categories of devices (robots, computers, guest devices, etc).
Our router and ad-hoc setups typically operate at a broadcast channel capacity of , and we typically observe around network load when operating 5 robots concurrently in a centralized mode (transmitting wheel velocity commands to each robot at 50 Hz). For details on our Ad-hoc peer-to-peer network, we refer the reader to [15] and [16].
@@
Crash prevention is a priority in high-velocity multi-robot experiments, and can be particularly complex to implement in a fully decentralized framework. Inevitable system latencies in re-establishing SSH sessions and the finite interval between terminating a control process and robot immobilization necessitate a more reliable stop mechanism. Consequently, we integrate an emergency stop service into our RoboMaster ROS driver, accessible via the infrastructure network, that immediately halts velocity commands at the lowest level. Additionally, we developed a stand-alone physical emergency stop button, powered by a Raspberry Pi Zero, that connects to the infrastructure network and features color LEDs to display the connectivity of the robots.
@@
Incorporating on-robot displays has proven advantageous, especially for operations outside the conventional laboratory setting where an infrastructure network may be absent. These displays, compact and economical at pixels, present vital statistics including CPU and memory usage, CPU temperature, network details (such as IP addresses and interface connectivity), battery status, and the robot’s hostname. An integrated button activates the emergency stop service if the robot is in motion. This feature is particularly useful in field deployments, offering a direct control and monitoring mechanism when network access is unavailable.
@@ 3.7 Simulation: VMAS
To enable this deployment pipeline, we integrated several new features into the simulator. Firstly, to generate the input required by Freyja, we implement a PID controller layer that regulates the input forces required by VMAS based on these reference velocities. This layer allows us to train neural network policies that output desired velocities and directly use them on our physical platform, without any change. Secondly, a key to the real-world deployment was an iterative fine-tuning process for some of the physics parameters in VMAS. These are parameters that depend on the robot platform and its interaction with the environment, such as friction coefficients, drag, control ranges, state boundaries, and robot geometry. This tuning process can be done once for each robot-environment pair and can be then utilized for any subsequent deployments.
A model for the RoboMaster platform in VMAS enables applications beyond prototyping in simulation and deploying in the real world. For instance, thanks to the accelerated computing available on our platform, policies trained in VMAS on approximated representations of the environment can be fine-tuned on the robot using data collected in the real world.
@@ 4 Evaluations
We first demonstrate the agility of our platform in a single-robot experiment in which we benchmark the tracking performance of Freyja as well as the velocity of our robot platform at high speeds. We demonstrate this on a straight line trajectory and a circle trajectory in Fig. 6. The straight line trajectory has a length of ca. , and the circle trajectory has a diameter of . We show that the robot reaches a top velocity of , thus exceeding the maximum rated speed of the DJI RoboMaster platform by , and accelerations of up to in the straight line trajectory with a maximum error of , and maintains a velocity of on the circle trajectory with an average error of . In the line trajectory, we are constrained by the size of the laboratory, which prohibits the robots from going faster.
The platform is capable of at least 600 deg/s in angular velocity, however, being a holonomic system, we do not stress this any further.
@@
for a multi-robot navigation task trained in VMAS [10] using the BenchMARL library [11] and TorchRL [17], following the pipeline described in Sec. 3.7.
Agents are rewarded for navigating to their goal while avoiding collisions with each other, to do so they use the GPPO model ([12]), which leverages a Graph Neural Network (GNN) in the policy for inter-agent communication. The GNN utilizes a 5-layer MLP with a total of 2300 trainable parameters, which can be evaluated on the Jetson Orin NX in . The results, reported in Fig. 7, show that eight agents are able to seamlessly execute the learned policy collision-free in the previously unseen real-world environment.
@@
a nonlinear model predictive controller to avoid collisions with both peers and static obstacles.
Fig. 8 tests a collision avoidance scenario at an intersection, where two robots (travelling along the horizontal and vertical axes) would nominally collide at the intersection.
However, they are able to localize each other close to the origin when common features are available in the environment, and the robot travelling along the vertical axis is able to slow down to avoid colliding. We are including additional details of this setup in a follow-up work.
Multi-robot control using CoViS-Net.
In this experiment, we deploy CoViS-Net [16], a pre-trained neural network to estimate relative poses between two camera frames, on two different robots. The camera images of two robots are encoded, and the features are broadcasted over the ad-hoc network to other robots within the communication range, based on which the relative pose can be predicted, which is then used to track a trajectory using Freyja.
@@
Case Studies. Our system has reached this level of maturity over the course of several iterations of research and development in the last four years. Apart from the capabilities demonstrated above, this framework has featured as a primary research platform in several prior multi-agent works from our group: